# Value Residual Learning

### v4

@@ Abstract
While Transformer models have achieved remarkable success in various domains, the effectiveness of information propagation through deep networks remains a critical challenge. Standard hidden state residuals often fail to adequately preserve initial token-level information in deeper layers. This paper introduces ResFormer, a novel architecture that enhances information flow by incorporating value residual connections in addition to hidden state residuals. And a variant is the SVFormer, where all layers share the first layer’s value embedding. Comprehensive empirical evidence demonstrates ResFormer achieves equivalent validation loss with 13.3% fewer model parameters and 15.4% less training data compared to Transformer, while maintaining similar memory usage and computational cost. Besides, SVFormer reduces KV cache size by nearly half with only a small performance penalty and can be integrated with other KV-efficient methods, yielding further reductions in KV cache, with performance influenced by sequence length and cumulative learning rate.
@@
Zhanchao Zhou1,2   Tianyi Wu3†thanks: Equal Contribution; †Work done during internship at Westlake University; $\diamond$ Corresponding author.   Zhiyun Jiang4†∗   Fares Obeid5   Zhenzhong Lan2⋄ 1Zhejiang University      2Westlake University 3University of Electronic Science and Technology of China 4China University of Mining and Technology     5Imperial College London
@@ 1 Introduction
In a standard decoder-only transformer, initial token embeddings contain localized information, which rapidly e volves into abstract semantic features through early attention layers (Sun et al., 2024b; Clark et al., 2019). As Transformer deepens, a critical question arises: How effectively is the initial information propagated to deeper layers? One common answer is that residual connections of hidden states ensure access to initial information throughout the network. However, some studies (Zhou et al., 2021; Shi et al., 2022) have identified that the smoothing effect of attention mechanisms leads to over-smoothing, where token representations become increasingly similar as the network deepens. This indicates that in deeper layers, sequence-level features become dominant, while token-level features are diluted. DenseFormer (Pagliardini et al., 2024) applied the idea of learnable dense connections from DenseNet (Huang et al., 2016) to transformers, and the learned connection coefficients shows that deeper layers indeed require larger attention to initial embeddings. Given the low similarity between initial token embeddings and deeper hidden states (Sun et al., 2024b), their directly summation may significantly impact the modeling of attention distribution for abstract semantic information in later layers. NeuTRENO (Nguyen et al., 2023) alleviates over-smoothing from the view of regularizers by considering the difference between value vectors of the first and current layers.
@@
During inference, deep networks require substantial $KV$ cache, severely impacting model deployment (Xiao et al., 2024). Existing $KV$-efficient methods often process keys and values simultaneously. Building on ResFormer, we decouple the value from the attention operation and propose a new kind of Transformer (SVFormer) where all layers share a single value states.
We experiment on a 20B SlimPajama sub-sampled dataset, using settings similar to popular large language models (Wei et al., 2023; Dubey et al., 2024; Kaplan et al., 2020). We compare different models based on the valid loss against the vanilla Transformer. Results show that ResFormer outperforms the vanilla Transformer, DenseFormer, and NeuTRENO. ResFormer achieves equivalent validation loss with 13.3% fewer model parameters and 15.4% less training data compared to Transformer, while maintaining similar memory usage and computational cost. Besides, SVFormer, while reducing the $KV$-cache by nearly half, requires a 12.2% increase in parameters to achieve the same validation loss as Transformer. And SVFormer performs better when the training sequence length is longer. It further reduces the $KV$ cache when integrated with GQA (Ainslie et al., 2023).
@@ 3 Method
3.1 Notations
Let $\mathbf{H}_{n}\in\mathbb{R}^{l\times d}$ be the input hidden state of the $n$-th layer, where $l$ denotes the sequence length and $d$ is the dimension size. For each layer, the hidden state $\mathbf{H}_{n-1}$ will be firstly projected into $\mathbf{Q}_{n},\mathbf{K}_{n},\mathbf{V}_{n}\in\mathbb{R}^{l\times d}$ through three linear projections $\mathbf{W^{\mathbf{Q}}_{n}},\mathbf{W^{\mathbf{K}}_{n}},\mathbf{W^{\mathbf{V}}_{n}}\in\mathbb{R}^{d\times d}$ respectively. After these projections, the attention operation ($\operatorname{Attn}$), output projection ($\mathbf{W}^{\mathbf{O}}_{n}\in\mathbb{R}^{d\times d}$), and Multi-Layer-Perceptron ($\operatorname{Mlp}$) are applied sequentially:
@@ 3.2 NeuTRENO and DenseFormer
$\mathbf{H}_{n}$ $=\bm{\lambda}_{n,n}\operatorname{Mlp}(\mathbf{U}_{n}\mathbf{W}^{\mathbf{O}}_{n})+\sum_{i=1}^{n-1}\bm{\lambda}_{n,i}\mathbf{H}_{i}.$ (4)
where $\mathbf{H}_{0}=\operatorname{Embedding}(\mathbf{X})$ for the input $\mathbf{X}$. $\bm{\lambda}_{n}$ in Eqn. 3 and $\!\{\bm{\lambda}_{n}\!\}$ in Eqn. 4 are new parameters.
@@ 3.3 ResFormer
In contrast, before Eqn. 1, ResFormer introduces a skip connection from the first layer’s value $\mathbf{V}_{0}=\mathbf{H}_{0}\mathbf{W}^{\mathbf{V}}_{1}$ to current layer’s value $\mathbf{V}_{n}=\mathbf{H}_{n}\mathbf{W}^{\mathbf{V}}$:
$\mathbf{V}_{n}$ $={\bm{\lambda}_{n,1}}\mathbf{V}_{1}+{\bm{\lambda}_{n,2}}\mathbf{H}_{n}\mathbf{W}^{\mathbf{V}}.$ (5)
@@
When all $\bm{\lambda}_{1}$ and $\bm{\lambda}_{2}$ are predetermined constants, it is termed Constant-ResFormer. If $\{{\bm{\lambda}_{n,1}}={\bm{\lambda}_{n,2}}\}_{n=1}^{N}$, where $N$ is the total number of layers, the model is called Identity-ResFormer. Another variant where some layers have $\bm{\lambda}_{1}=0$ is referred to as Sparse-ResFormer. Besides, if $\bm{\lambda}_{1}$ and $\bm{\lambda}_{2}$ are trainable parameters, the model is termed Learnable-ResFormer. Unless otherwise specified, $\bm{\lambda}_{1}$ and $\bm{\lambda}_{2}$ are initialized to 0.5 for Learnable-ResFormer and are predetermined as 0.5 for Identity-ResFormer.
A more general form is the Dense-ResFormer, defined as $\mathbf{V}_{n}\!=\!\bm{\lambda}_{n,n}\mathbf{H}_{n}\mathbf{W}^{\mathbf{V}}+\sum_{i=1}^{n-1}\bm{\lambda}_{n,i}\mathbf{V}_{i}$ for $n\geq 2$, where $\{\bm{\lambda}_{n}\!\}$ are constants or trainable scalars. Unless noted, all $\bm{\lambda}$ are initialized to 1 here.
@@ 3.4 SVFormer
Beyond ResFormer, SVFormer adopts standard attention in the first layer and obtain the attention output $\mathbf{U_{n}}$ for $n$-th layer where $n\geq 2$ through $\mathbf{U}_{n}=\mathbf{A}_{n}{\mathbf{V}_{1}}.$ Its main advantage is that it only requires computing and storing the value vectors for the first layer, saving nearly half of the $KV$ cache during inference. Similar methods like CLA reduce $KV$ cache by sharing both of the key and value vectors every two layers. However, the results in Table  1 show that sharing values has less negative impact compared with sharing keys.
@@ 4.1 Setting
Following Brandon et al. (2024), we choose the Llama-like architecture and SlimPajama (Soboleva et al., 2023) data for main experiments. Specifically, the architecture includes pre-normalization, SwiGLU activations (Shazeer, 2020), rotary position embedding (Su et al., 2024), and no dropout. For slimpajama, we randomly sample nearly 20B tokens based to the original data distribution of seven domains during training and adopt tokenizer used for “RedPajama-INCITE-7B-Base". See Table 12 in Appendix for data details.
@@
4.2 Scaling Analysis of ResFormer
Figure 2: (Left) Validation loss as model size scales from 82M to 468M parameters. (Right) Validation loss for the 468M parameter model evaluated every 2B tokens. ResFormer achieves approximately 13.3%-15.4% reduction in both model parameters and training data.
We analyzed how ResFormer and Transformer scale at model size and data size. ResFormer and Transformer are trained on similar experiment setting. On the one hand, we trained model with 82M, 180M, 320M and 468M parameters on 20B training tokens and evaluated them on a separate validation set. As shown in Fig.2 (Left), ResFormer achieves equivalent validation loss to the Transformer while utilizing 13.3% fewer model parameters. On the other hand, we evaluated the 468M models every 2B tokens and ResFormer needs 15.4% fewer training tokens to achieve the same loss as Transformer.
@@ 4.2 Scaling Analysis of ResFormer
$\text{DenseFormer}^{*}$ $\mathbf{H}_{n}^{\prime}\!=\!{\color[rgb]{1,0,0}1\times}\mathbf{H}_{i}+\sum_{i=1}^{n-1}{\color[rgb]{1,0,0}0\times}\mathbf{H}_{i}$ 2.722
@@
Constant-ResFormer $\mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}2}\mathbf{V}_{1}+{\color[rgb]{1,0,0}0.5}\mathbf{V}_{n}$ 2.7
@@
Sparse-ResFormer $\left\{\begin{array}[]{l}\mathbf{V}_{n}^{\prime}=\mathbf{V}_{n},1\leq n\leq 5\\ \mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}5}\mathbf{V}_{1}+{\color[rgb]{1,0,0}0.5}\mathbf{V}_{n},\\ \qquad\ \quad\qquad 6\leq n\leq 8\end{array}\right.$ 2.687
@@
We compared the average validation loss of different models with 82M parameters trained on 20B tokens. In our experimental setup, NeuTRENO performed best with $\bm{\lambda}=0.4$, achieving comparable results to DenseFormer. All ResFormer variants, including the simplest Identity-ResFormer, demonstrated significant performance improvements. The $\bm{\lambda}$ values for Constant-ResFormer and Sparse-ResFormer were optimized through multiple experiments. While Sparse-ResFormer achieved the best overall performance, it is challenging to pre-determine the optimal layers for $\mathbf{V}_{1}$ connections and their corresponding $\bm{\lambda}_{1}$ values in more general scenarios. Although, Learnable-ResFormers are struggle to identify the optimal value residual pattern, they outperform Identity-ResFormer. Interestingly, the second to last row shows that Sparse-ResFormer achieved better performance despite having three fewer $\mathbf{W}^{\mathbf{V}}$.
@@ 4.5 Ablation Study of Value Residual
value residual $\mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}0.5}\mathbf{V}_{1}+{\color[rgb]{1,0,0}0.5}\mathbf{V}_{n}$ 2.7049
hidden residual $\mathbf{H}_{n}^{\prime}={\color[rgb]{1,0,0}0.5}\mathbf{H}_{0}+{\color[rgb]{1,0,0}0.5}\mathbf{H}_{n}$ 2.7812
value residual $\mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}0\times}\mathbf{V}_{1}+{\color[rgb]{1,0,0}1}\mathbf{V}_{n}$ 2.7298
hidden residual $\mathbf{H}_{n}^{\prime}={\color[rgb]{1,0,0}0\times}\mathbf{H}_{0}+{\color[rgb]{1,0,0}1}\mathbf{H}_{n}$ 2.7216
@@
Our experiments revealed that $\mathbf{V}_{1}$ information provides additional benefits to later network layers, despite both $\mathbf{H}_{1}$ and $\mathbf{V}_{1}$ containing initial, unfused token information. $\mathbf{H}_{0}$ is propagated through default hidden residual connections, but it may be diluted by subsequent information, hindering its effective utilization in later layers. To test this hypothesis, we introduced an additional skip connection to $\mathbf{H}_{0}$ : $\mathbf{H}_{n}^{\prime}={\mathbf{\lambda}_{1}}\mathbf{H}_{1}+{\mathbf{\lambda}_{2}}\mathbf{H}_{n}$, where $\mathbf{\lambda}$ is learnable. We conducted experiments with two $\mathbf{\lambda}$ initialization settings and compared them to value residual.
@@
$\mathbf{H_{0}}$ (Default) - 2.7389
$\mathbf{H_{0}}$ (Default) $\mathbf{V_{1}}$ 2.7117
$\mathbf{H_{0}}$ (Default) $\mathbf{V_{2}}$ 2.7375
$\mathbf{H_{1}}$ 2.7802
$\mathbf{H_{1}}$ $\mathbf{V_{2}}$ 2.784
$\mathbf{H_{2}}$ 2.8196
$\mathbf{H_{2}}$ $\mathbf{V_{2}}$ 2.7873
$\mathbf{H_{2}}$ 2.8196
$\mathbf{H_{2}}$ $\mathbf{V_{3}}$ 2.8333
$\mathbf{H_{3}}$ 3.0568
$\mathbf{H_{3}}$ $\mathbf{V_{3}}$ 2.8832

### v5

@@ Abstract
While Transformer models have achieved remarkable success in various domains, the effectiveness of information propagation through deep networks remains a critical challenge. Standard hidden state residuals often fail to adequately preserve initial token-level information in deeper layers. This paper introduces ResFormer, a novel architecture that enhances information flow by incorporating value residual connections in addition to hidden state residuals. And a variant is SVFormer, where all layers share the first layer’s value embedding. Comprehensive empirical evidence demonstrates ResFormer achieves equivalent validation loss with 16.11% fewer model parameters and 20.3% less training data compared to Transformer, while maintaining similar memory usage and computational cost. Besides, SVFormer reduces KV cache size by nearly half with only a small performance penalty and can be integrated with other KV-efficient methods, yielding further reductions in KV cache, with performance influenced by sequence length and cumulative learning rate.
@@
Zhanchao Zhou1,2,3   Tianyi Wu4†thanks: Equal Contribution; †Work done during internship at Westlake University; $\diamond$ Corresponding author.   Zhiyun Jiang5†∗   Fares Obeid6   Zhenzhong Lan2⋄ 1Zhejiang University      2Westlake University      3Ant Group 4University of Electronic Science and Technology of China 5China University of Mining and Technology     6Imperial College London
@@ 1 Introduction
In a standard decoder-only transformer, initial token embeddings contain localized information, which rapidly evolves into abstract semantic features through early attention layers (Sun et al., 2024b; Clark et al., 2019). As Transformers deepen, a critical question arises: How effectively is the initial information propagated to deeper layers? One common answer is that residual connections of hidden states ensure access to initial information throughout the network. However, some studies (Zhou et al., 2021; Shi et al., 2022) have identified that the smoothing effect of attention mechanisms leads to over-smoothing, where token representations become increasingly similar as the network deepens. This indicates that in deeper layers, sequence-level features become dominant, while token-level features are diluted. DenseFormer (Pagliardini et al., 2024) applied the idea of learnable dense connections from DenseNet (Huang et al., 2016) to Transformer, and the learned connection coefficients shows that deeper layers indeed require larger attention to initial embeddings. Given the low similarity between initial token embeddings and deeper hidden states (Sun et al., 2024b), their directly summation may significantly impact the modeling of attention distribution for abstract semantic information in later layers. NeuTRENO (Nguyen et al., 2023) alleviates over-smoothing from the view of regularizers by considering the difference between value vectors of the first and current layers.
@@
During inference, deep networks require substantial $KV$ cache, severely impacting model deployment (Xiao et al., 2024). Existing $KV$-efficient methods often process keys and values simultaneously. Building on ResFormer, we decouple the value from the attention operation and propose a new kind of Transformer (SVFormer) where all layers share a single value state.
We experiment on a 20B SlimPajama sub-sampled dataset, using settings similar to popular large language models (Wei et al., 2023; Dubey et al., 2024; Kaplan et al., 2020). We compare different models based on the valid loss against the vanilla Transformer. Results show that ResFormer outperforms the vanilla Transformer, DenseFormer, and NeuTRENO. ResFormer achieves equivalent validation loss with 16.11% fewer model parameters and 20.3% less training data compared to Transformer, while maintaining similar memory usage and computational cost. Besides, SVFormer, while reducing the $KV$-cache by nearly half, requires a 12.2% increase in parameters to achieve the same validation loss as Transformer. And SVFormer performs better when the training sequence length is longer. It further reduces the $KV$ cache when integrated with GQA (Ainslie et al., 2023).
@@ 3 Method
3.1 Preliminary
Notations
Let $\mathbf{H}_{n}\in\mathbb{R}^{l\times d}$ be the output hidden state of the $n$-th layer, where $l$ denotes the sequence length and $d$ is the dimension size. For each layer, the hidden state $\mathbf{H}_{n-1}$ will be firstly projected into $\mathbf{Q}_{n},\mathbf{K}_{n},\mathbf{V}_{n}\in\mathbb{R}^{l\times d}$ through three linear projections $\mathbf{W^{\mathbf{Q}}_{n}},\mathbf{W^{\mathbf{K}}_{n}},\mathbf{W^{\mathbf{V}}_{n}}\in\mathbb{R}^{d\times d}$ respectively. After these projections, the attention operation ($\operatorname{Attn}$), output projection ($\mathbf{W}^{\mathbf{O}}_{n}\in\mathbb{R}^{d\times d}$), and Multi-Layer-Perceptron ($\operatorname{Mlp}$) are applied sequentially:
@@ 3.1 Preliminary
$\mathbf{H}_{n}$ $=\bm{\lambda}_{n,n}\operatorname{Mlp}(\mathbf{U}_{n}\mathbf{W}^{\mathbf{O}}_{n})+\sum_{i=0}^{n-1}\bm{\lambda}_{n,i}\mathbf{H}_{i}.$ (4)
where $\mathbf{H}_{0}=\operatorname{Embedding}(\mathbf{X})$ for the input $\mathbf{X}$. $\bm{\lambda}_{n}$ in Eqn. 3 and $\!\{\bm{\lambda}_{n,i}\!\}_{i=0}^{n-1}$ in Eqn. 4 are new parameters. Unless noted, $\bm{\lambda}_{n}$ is set to 0.4 for NeuTRENO suggested by Nguyen et al. (2023). For DenseFormer, only $\bm{\lambda}_{n,n}$ are set to 1 and the others are set to zero during initialization here.
@@ 3.2 ResFormer
In contrast, before Eqn. 1, ResFormer introduces a skip connection from the first layer’s value $\mathbf{V}_{1}=\mathbf{H}_{0}\mathbf{W}^{\mathbf{V}}_{1}$ to current layer’s value $\mathbf{V}_{n}=\mathbf{H}_{n-1}\mathbf{W}^{\mathbf{V}}_{n}$:
$\mathbf{V}_{n}$ $={\bm{\lambda}_{n,1}}\mathbf{V}_{1}+{\bm{\lambda}_{n,2}}\mathbf{H}_{n-1}\mathbf{W}^{\mathbf{V}}_{n}.$ (5)
@@
When all $\bm{\lambda}_{n,1}$ and $\bm{\lambda}_{n,2}$ are predetermined constants, it is termed Constant-ResFormer. If $\{{\bm{\lambda}_{n,1}}={\bm{\lambda}_{n,2}}\}_{n=1}^{N}$, where $N$ is the total number of layers, the model is called Identity-ResFormer. Another variant where some layers have $\bm{\lambda}_{n,1}=0$ is referred to as Sparse-ResFormer. Besides, if $\bm{\lambda}_{n,1}$ and $\bm{\lambda}_{n,2}$ are trainable parameters, the model is termed Learnable-ResFormer. Unless otherwise specified, $\bm{\lambda}_{n,1}$ and $\bm{\lambda}_{n,2}$ are initialized to 0.5 for Learnable-ResFormer and are predetermined as 0.5 for Identity-ResFormer. Furthermore, given that higher layers require more supplementary information from $\mathbf{V}_{1}$, we propose the Learnable ResFormer Plus. The initialization strategy is as follows: 1). $\bm{\lambda}_{n,2}$ is initialized to 0.5, where $n=1,2,\ldots,N$. 2). $\bm{\lambda}_{n,1}$ is initialized to $\bm{\lambda_{\text{scale}}}\cdot\frac{e^{\bm{\lambda}_{n,1}^{\prime}}}{\sum_{j=1}^{N}e^{\bm{\lambda}_{j,1}^{\prime}}}$, where $\bm{\lambda_{\text{scale}}}$ is initialized to the total layer number $N$ and is shared across all layers.
A more general form is Dense-ResFormer, defined as $\mathbf{V}_{n}\!=\!\bm{\lambda}_{n,n}\mathbf{H}_{n-1}\mathbf{W}^{\mathbf{V}}_{n}+\sum_{i=1}^{n-1}\bm{\lambda}_{n,i}\mathbf{V}_{i}$ for $n\geq 2$, where $\!\{\bm{\lambda}_{n,i}\!\}_{i=1}^{n-1}$ are constants or trainable scalars. Unless noted, all $\bm{\lambda}_{n,n}$ are set to 1.
@@ 3.3 SVFormer
Figure 2: (Left) Validation loss as model size scales from 82M to 468M parameters on 20B tokens. (Medium) Validation loss for the 468M parameter model evaluated every 2B tokens. ResFormer achieves approximately 16.1%-20.3% reduction in both model parameters and training data. (Right) Validation loss for the 1.6B parameter model evaluated every 10B tokens.
Model Wiki. LMB. Arc-c Arc-e BoolQ Hella. LMB. Obqa. Wino. PiQA Avg.
PPL PPL ACC ACC ACC ACC ACC ACC ACC ACC ACC
Transformer 24.8 33.4 18.8 49.6 57.5 31.1 34.1 17.6 49.8 66.1 40.6
NeuTRENO 24.3 36.3 20.7 48.8 59.2 31.6 34.2 19.0 51.8 65.9 41.4
DenseFormer 24.0 28.0 20.1 48.7 56.5 32.1 36.6 17.6 49.3 65.6 40.8
Identity ResFormer 23.8 32.9 20.2 49.0 59.0 32.1 36.0 16.8 51.2 66.1 41.3
Learnable ResFormer 23.7 32.5 21.2 50.3 60.9 32.3 36.3 18.8 51.2 67.0 42.3
Learnable ResFormer plus 23.2 31.4 21.2 49.7 60.6 32.4 36.0 17.8 51.1 67.5 42.0
Table 2: Downstream evaluation of different models with 468M parameters trained on 20B tokens.
Beyond ResFormer, SVFormer adopts standard attention in the first layer and obtain the attention output $\mathbf{U_{n}}$ for $n$-th layer when $n\geq 2$ through $\mathbf{U}_{n}=\mathbf{A}_{n}{\mathbf{V}_{1}}$, where $\mathbf{A}_{n}$ is the attention matrix of $n$-th layer. Its main advantage is that it only requires computing and storing the value vectors for the first layer, saving nearly half of the $KV$ cache during inference. Similar methods like CLA reduce $KV$ cache by sharing both of the key and value vectors every two layers. However, the results in Table  1 show that sharing values has less negative impact compared with sharing keys.
@@ 4.1 Setting
Following Brandon et al. (2024), we choose the Llama-like architecture and SlimPajama (Soboleva et al., 2023) data for main experiments. Specifically, the architecture includes pre-normalization, SwiGLU activations (Shazeer, 2020), rotary position embedding (Su et al., 2024), and no dropout. For SlimPajama, we randomly sample nearly 20B tokens based on the original data distribution of seven domains during training and adopt tokenizer used for “RedPajama-INCITE-7B-Base". See Table 12 in Appendix for data details.
@@
Scaling Details
We conduct scaling experiments on a 1.6B parameter model using approximately 200B tokens of internal pre-training data. Training is performed with a sequence length of 8,192, learning rate of 1.5e-4, and batch size of 2M tokens, requiring about 3,584 H800 GPU hours per run.
@@
We also compare different models on several classical reasoning tasks following (Zhang et al., 2024) in a zero-shot way. The tasks include Hellaswag (Hella.) (Zellers et al., 2019) Openbookqa (Obqa.) (Mihaylov et al., 2018), WinoGrande (Wino.) (Sakaguchi et al., 2019), ARC-Easy (Arc-e) and ARC-Challenge (Arc-c) (Clark et al., 2018) and PiQA (Bisk et al., 2020). In addition to accuracy on reasoning tasks, we also report the perplexity (PPL) on Wikitext (Wiki.) and LAMBADA (LMB.).
@@
We analyze how different models scale with model size and data size under similar experimental settings. We train models with 82M, 180M, 320M, and 468M parameters on 20B tokens and evaluate on a validation set. Fig.2 (Left) shows ResFormer achieves equivalent validation loss to Transformer while using 16.11% fewer parameters. We also evaluate 468M models every 2B tokens, finding ResFormer requires 20.3% fewer training tokens to match Transformer’s loss (Fig.2 (Medium)). All ResFormer variants demonstrate superior scaling compared to NeuTRENO and DenseFormer. Table 2 confirms ResFormer outperforms other models on downstream tasks, with Learnable ResFormer achieving 1.7 point average accuracy improvement over vanilla Transformer.
To validate the performance improvement of value residual at larger training scales, we conduct additional experiments following the scaling details in Sec.4.1. As shown in Fig.2 (Right), value residual consistently improves performance throughout training.
@@
$\text{DenseFormer}^{*}$ $\mathbf{H}_{n}^{\prime}\!=\!{\color[rgb]{1,0,0}1\times}\mathbf{H}_{n}+\sum_{i=0}^{n-1}{\color[rgb]{1,0,0}0\times}\mathbf{H}_{i}$ 2.722
@@
Constant-ResFormer $\mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}2}\mathbf{V}_{1}+\mathbf{V}_{n}$ 2.7
@@
Sparse-ResFormer $\left\{\begin{array}[]{l}\mathbf{V}_{n}^{\prime}=\mathbf{V}_{n},1\leq n\leq 5\\ \mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}5}\mathbf{V}_{1}+\mathbf{V}_{n},\\ \qquad\ \quad\qquad 6\leq n\leq 8\end{array}\right.$ 2.687
[…]
@@
Furthermore, we test the performance of other varaints of ResFormer mentioned in Sec.3.2. The $\bm{\lambda}$ values for Constant-ResFormer and Sparse-ResFormer were optimized through multiple experiments. All ResFormer variants, including the simplest Identity-ResFormer, show significant performance improvements. However, manually-tuned Sparse-ResFormer and Learnable-ResFormer-Plus outperform the standard Learnable-ResFormer. It demonstrates the challenge to pre-determine the optimal layers for $\mathbf{V}_{1}$ connections and their corresponding $\bm{\lambda}_{1}$ values in more general scenarios. Interestingly, the third to last row shows that Sparse-ResFormer achieved better performance despite having three fewer $\mathbf{W}^{\mathbf{V}}$.
@@ 4.4 Ablation Study of Value Residual
value $\mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}0.5}\mathbf{V}_{1}+{\color[rgb]{1,0,0}0.5}\mathbf{V}_{n}$ 2.705
hidden $\mathbf{H}_{n}^{\prime}={\color[rgb]{1,0,0}0.5}\mathbf{H}_{0}+{\color[rgb]{1,0,0}0.5}\mathbf{H}_{n}$ 2.781
value $\mathbf{V}_{n}^{\prime}={\color[rgb]{1,0,0}0\times}\mathbf{V}_{1}+{\color[rgb]{1,0,0}1}\mathbf{V}_{n}$ 2.73
hidden $\mathbf{H}_{n}^{\prime}={\color[rgb]{1,0,0}0\times}\mathbf{H}_{0}+{\color[rgb]{1,0,0}1}\mathbf{H}_{n}$ 2.722
@@
Our experiments revealed that $\mathbf{V}_{1}$ information provides additional benefits to later network layers, despite both $\mathbf{H}_{0}$ and $\mathbf{V}_{1}$ containing initial, unfused token information. $\mathbf{H}_{0}$ is propagated through default hidden residual connections, but it may be diluted by subsequent information, hindering its effective utilization in later layers. To test this hypothesis, we introduced an additional skip connection to $\mathbf{H}_{0}$ : $\mathbf{H}_{n}^{\prime}={\mathbf{\lambda}_{0}}\mathbf{H}_{1}+{\mathbf{\lambda}_{2}}\mathbf{H}_{n}$, where $\mathbf{\lambda}$ is learnable. We conducted experiments with two $\mathbf{\lambda}$ initialization settings and compared them to value residual.
@@
$\mathbf{H_{0}}$ (Default) - 2.739
$\mathbf{H_{0}}$ (Default) $\mathbf{V_{1}}$ 2.712
$\mathbf{H_{0}}$ (Default) $\mathbf{V_{2}}$ 2.738
$\mathbf{H_{1}}$ 2.78
$\mathbf{H_{1}}$ $\mathbf{V_{2}}$ 2.78
$\mathbf{H_{2}}$ 2.82
$\mathbf{H_{2}}$ $\mathbf{V_{2}}$ 2.787
$\mathbf{H_{2}}$ 2.82
$\mathbf{H_{2}}$ $\mathbf{V_{3}}$ 2.833
$\mathbf{H_{3}}$ 3.057
$\mathbf{H_{3}}$ $\mathbf{V_{3}}$ 2.883